Aug 6, 2009 Calorimetry

From GlueXWiki
Jump to: navigation, search

Teleconference Time: 11:00 a.m. EST

  • ESNET and EVO session
  • Phone:
    • +1-800-377-8846 : US
    • +1-888-276-7715 : Canada
    • +1-302-709-8424 : International
    • then enter participant code: 77438230# (remember the "#"). Note: code changed 7/2/09

Items for followup from previous meeting(s)

  1. Procedure to avoid vaseline in Bcal fabrication needs to be approved.
  2. Documentation on fiber first-article
  3. Review calibration procedures

Documents to Review

Tentative Agenda

  1. Announcements
  2. Review of Action Items
  3. FCAL Update
  4. BCAL Updates
    1. Fabrication (Zisis)
      1. Press Frame
      2. Swager/Lead
      3. Fibre Sorting Table
      4. Construction Methods
    2. Fibre QA (Zisis)
  5. Simulations
  6. BCAL Readout
  7. Other business


Attendees: Jlab: Elton, Eugene, Beni, Fernando, Tim, Sasha, Ashot; USM: Alam, Will (traveling to NY); IU: Matt, Ryan; UofR: Zisis, George, Andrei

  1. Announcements
    • None
  2. Review of Action Items
    • First to items addressed. Dedicated calibration and monitoring meeting will be scheduled.
  3. FCAL Update
    1. CAN-ethernet bridge is working well. Test for use with slow controls by Claire was successful. Will be documented
    2. Contract
      1. PR for production of Fcal modules has been submitted and signed.
      2. Detailed cost response needs to be prepared. (Elton to send Regina docs to Matt for reference).
    3. FEU-84-3 pmt specs
      1. Matt will meet with Sergey Denisov next week. Needs to get the following information to enable the procurement to go forward:
        • contact info (person as well as e-mail/address/etc)
        • estimated cost
        • updates to specs as suggested by Sergey. Add "not-used" to specification.
        • How much (if any testing) can be done before shipping pmts? Should one plan to purchase more than needed and make a selection here? What fraction of rejects are expected?
    4. Cosmic-ray test setup -constant progress. Hope to have something to show at collaboration meeting
    5. Radiation damage tests (Ryan)
      1. Irradiated glass blocks at cyclotron.
      2. Brad (student) measuring transmission of irradiated glass
      3. Transmission vs wavelength for various doses has been completed.
      4. Still need to understand how to convert of registered dose to units such as Gray.
    6. Magnetic shields (Tim)
      1. 14 mil shield has been modified to fit inside shield -> send back to Amuneal to anneal shield and ship back
      2. Expect receipt of annealed shield in two weeks to verify final configuration.
  4. BCAL Updates
    1. Fabrication (Zisis)
      1. Press Frame
        • significant retrofits to new press frame; lunar lander feet installed; guide system to register to grooves;
      2. Dan's office setup next to production
      3. Swager/Lead setup into permanent location
        • 70% "Beaufait" lead used for prototype, 30% of first-article
        • Pitch of swager setup for Beaufait lead at 1.25 mm pitch. The design goal is 1.18 mm. Will readjust swager to first article lead and see if it comes close to design specs.
        • Creating a spare-parts list for swager machine.
      4. Lead
        • First article acceptable, although minor differences in thickness found between 10cm-wide and 11cm-wide strips. Difference vanishes after swaging.
        • (Tim) Expect shipment of first 25% of lead in 8 weeks. Smaller shipments could be made sooner if required.
      5. Fibre Sorting Table
        • fitted with new casters to allow ease of movement during full production
      6. Construction Methods
        • Doing de-lamination tests with various setups. Found a NO-vaseline solution with a polyethylene sheet + 3/4" silicon rubber.
    2. Contract
      1. Claus indicated that DOE is close to signing contract
      2. Elton has collected the relevant information for a CR (change request) to the official schedule based on the Bcal fabrication contract. This will facilitate reporting on the "earned value" for the Bcal construction.
    3. Fibre QA (Zisis)
      1. Shipment03 Preliminary JPG Media:Shipment03_Preliminary.jpg
      2. Shipment03 Preliminary PDF Media:Shipment03_Preliminary.pdf
      3. Measured about 50 fibers. Data show consistent results compared to previous shipments, meeting specifications. There are still some fibers that showed some inconsistencies and will be retested.
    4. First-article draft in circulation between UofR and Jlab.
      1. Includes UofR reanalysis of JLab data using UofR extraction of Npe.
      2. Will be posted to the portal once comments are included.
  5. Simulations
    1. (Andrei)
      • Checked the simulated pulses in the summed readout output for 60 and 90 MeV (90 deg). These were showing about 15% zeros (no signal).
      • This was investigated and is believed to be due to physical effects (not bugs in code). Checked that there are non-zero signals in successive readout channels (4+5+6).
      • The conclusion is that indeed summing of signals into readout channels increases the chance that signals will be measured that are above threshold.
  6. BCAL Readout
    1. No meeting next week.
  7. Other business
    1. Discussed focus meeting on calibration and monitoring